# Rising bubbles draw surface patterns: a numerical study

### v1

@@ III.3 Multi-armed modes
Figure 10: (a) Snapshot of a cross-sectional view of the bubbles near the bump. The streamlines are colored according to the velocity magnitude. The dashed line indicates the central axis. (b) Contour lines of the liquid’s radial velocity in a horizontal plane slightly below the free surface. The small circles are bubbles floating on the surface.
@@
After escaping from the bump, in stage VI, bubble 0 assumes a purely radial motion, as evidenced by the negligible circumferential velocity shown in Fig. 9(b). According to the contour lines in Fig. 10(b), the radial velocity of the bubbles is smaller than that of the surrounding liquid, revealing that the bubbles are passively advected by the radial liquid flow in the surface layer. During this stage, both the radial and vertical velocity components of the bubble undergo mild fluctuations over time, as shown in Fig. 9(b). These fluctuations are likely caused by the outward-propagating annular surface waves faintly visible in the top row of Fig. 4. Due to volume conservation and viscous dissipation, the radial flow weakens with increasing radius. The bubbles gradually decelerate and eventually stop at some distance from the center.
IV Further theoretical considerations
@@
We propose a simple model to describe the dependence of the divergence angle $\varphi$ on the bubble release period $T$, provided that all other parameters are fixed. Without loss of generality, we assume the motion of bubble 2-2 is in the direction of x-x, as illustrated in Fig. 11(a). Based on the discussions in section III.3, in stages I and II, the motion of bubble 0 is mainly influenced by bubbles 2-2 and 1-1, with the displacements occurring in the opposite directions of these two bubbles. Therefore, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are plotted in Fig. 11(a) parallel to the reverse extensions of the two bubbles’ position vectors. In later stages, bubble 0 is launched in the direction of the vector sum of $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$. In order that the successive bubbles form a regular pattern, the angles between bubbles 0 and 1-1 and between bubbles 1-1 and 2-2 are both $\varphi$. With introduction of the variable $\alpha=\varphi-\pi/2$, one obtains $\beta=\pi/2-2\alpha$, where $\beta$ is defined in Fig. 11(a). Denoting the magnitude of $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ by $\lambda_{1}$ and $\lambda_{2}$, respectively, the geometric constraint requires $\tan\beta=(\lambda_{1}-\lambda_{2}\sin\alpha)/(\lambda_{2}\cos\alpha)$, which yields
@@
When $\lambda_{2}/\lambda_{1}\rightarrow 2^{-}$, $\varphi\rightarrow\pi$, corresponding to the two-armed mode limit. When $\lambda_{2}/\lambda_{1}\rightarrow 0^{+}$, $\varphi\rightarrow\pi/2$, indicating a four-armed mode. However, this latter situation is unlikely to occur because of the fact that while $\lambda_{2}$ is finite in general, $\lambda_{1}$, caused by non-contact repulsion, cannot be excessively large. The situation of $\lambda_{2}/\lambda_{1}>2$ is not covered by Eq. (6), either can it be described by the geometrical configuration shown in Fig. 11(a). Actually, it represents the scenario that the non-contact repulsion between bubbles 2-2 and 0 is too small to influence the pattern formation, and the divergence angle is dominated by the pair collision between bubbles 1-1 and 0. Thus, the fixed two-armed mode or disorder mode is expected.
Figure 11: (a) Geometric configuration of bubbles on the free surface (bubble size and position for illustrative purposes only). (b) Simulated bubble trajectories projected onto the horizontal plane with appropriate translation and rotation.
@@
Eq. (7) is valid unless $T$ is too large, in which case disordered, instead of regular, patterns, will be observed. Fig. 12(a) shows the averaged $\varphi$ as a function of $T$ for eight simulation cases (black circles) and the experimental data from Ref. [26] (red squares). Notably, with $a$ and $b$ fitted for each data set, the predictions of Eq. (7) (solid lines) have a remarkable consistency with both the numerical and experimental results. The fitted coefficients for the two curves, from left to right, are $a=0.039$ ms-1, $b=0.83$, and $a=0.035$ ms-1, $b=0.90$, respectively.
Figure 12: (a) Divergence angle $\varphi$ versus release period $T$. Red squares denote experimental data extracted from Fig. 1.2(a) in Ref. [26], with $Oh\approx 0.099$ and $Bo=17$. (b) Divergence angle versus dimensionless frequency $\Omega$. The three data sets on the left are experimental results extracted from Fig. 5(a) in Ref. [24]. From left to right, the Ohnesorge and Bond numbers, estimated based on the available information in Ref. [24], are approximately 0.26 and 23, 0.21 and 18, 0.11 and $1$, respectively. In both panels, the present numerical results are denoted by black circles; solid curves represent predictions using Eq. (7) or (8).
@@
where $\Omega=\rho_{l}d^{2}/(\mu_{l}T)$ is the dimensionless bubble release frequency defined in Ref. [24]. Correspondingly, $A=\rho_{l}d^{2}a/\mu_{l}$. The dependence of $\varphi/(2\pi)$ on $\Omega$ is shown in Fig. 12(b). The present numerical results are still denoted by black circles. The other three data sets correspond to experimental data extracted from Fig. 5(a) in Ref. [24]. The results of the theoretical model Eq. (8) are shown as black curves with $A$ and $b$ fitted for each set of data. The values of $A$ and $b$ for the four curves in Fig. 12(b), from left to right, are: 13.3 and 0.87, 16.5 and 0.64, 17.9 and 0.35, 31.1 and 0.83, respectively. Good agreement is observed between the theoretical predictions and all the data sets.
In the following, we consider the possibility of observing different multi-armed patterns in reality. Fig. 13(a) shows the number $n$ of arms that is expected to be seen as the release period $T$ varies. This mapping is obtained by calculating the corresponding $\varphi$ at a given $T$ using Eq. (7), then finding $n$ within the range $2\leq n\leq 15$ that best fits Eq. (5) with all possible $m$. The upper bound of 15 is set considering that for very large $n$, the finite-sized bubbles are so crowded that the arms become difficult to distinguish in practice. It can be seen from Fig. 13(a) that each $n$ may correspond to several discontinuous intervals of $T$. The accumulated range of $T$ for a given $n$ is shown in Fig. 13(b). The probability of appearances of $n$-armed mode can thus be estimated based on the height of the bars.
@@ V Summary and conclusions
This work elucidates the formation mechanism of regular surface patterns induced by chains of small rising bubbles. Direct numerical simulations reveal how distinct morphologies stem from non-contact bubble repulsion, pair collisions, free-surface deformation, and ambient flow advection. We establish a theoretical model relating the number of pattern arms to the bubble release period, showing quantitative agreement with both simulation data and experimental results from the literature. Future studies should explore other parameters including liquid viscosity, surface tension coefficient, and bubble diameter.
@@ ACKNOWLEDGEMENTS
D. Li and G. Zhou acknowledge the support of the National Natural Science Foundation of China (grant number 12202441). D. Li appreciates the advice from Dr. Petr Karnakov on the use of Aphros. The authors also thank Dr. Zibo Ren for his participation in the early stage of this study.

### v2

@@ III.3 Multi-armed modes
Figure 10: Snapshot of a cross-sectional view of the bubbles near the bump, with streamlines colored according to the velocity magnitude. The free surface is represented by the black solid line. The dashed line indicates the central axis.
Figure 11: Time dependence of bubble 0’s radial (blue), vertical (green), and circumferential (pink, clockwise positive) velocities. The red line represents the distance between the bubble mass center and the undisturbed free surface. Time is set to zero at the instant of bubble release.
@@
After escaping from the bump, in stage VI, bubble 0 assumes a purely radial motion. Figure 12 presents the contour lines of the liquid’s radial velocity in a horizontal plane slightly below the free surface. According to this figure, the radial velocity of the bubbles is smaller than that of the surrounding liquid, revealing that the bubbles are passively advected by the divergent liquid flow in the surface layer. During this stage, both the radial and vertical velocity components of the bubble undergo mild fluctuations over time, as can be observed in Fig. 11. These fluctuations are likely caused by the outward-propagating annular surface waves faintly visible in the top row of Fig. 4. Due to volume conservation and viscous dissipation, the radial flow weakens with increasing radius. The bubbles gradually decelerate and eventually stop at some distance from the center, forming discrete clusters, as is evident in Figs. 1, 4(c), and 4(d).
Figure 12: Contour lines of the liquid’s radial velocity in a horizontal plane slightly below the free surface. The small circles are bubbles floating on the surface.
IV A heuristic model
@@
We propose a simple model to describe the dependence of the divergence angle $\varphi$ on the bubble release period $T$, provided that all other parameters are fixed. Without loss of generality, we assume the motion of bubble 2-2 is in the direction of x-x, as illustrated in Fig. 13(a). Based on the discussions in section III.3, in stages I and II, the motion of bubble 0 is mainly influenced by bubbles 2-2 and 1-1, with the displacements occurring in the opposite directions of these two bubbles. Therefore, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are plotted in Fig. 13(a) parallel to the reverse extensions of the two bubbles’ position vectors. In later stages, bubble 0 is launched in the direction of the vector sum of $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$. In order that the successive bubbles form a regular pattern, the angles between bubbles 0 and 1-1 and between bubbles 1-1 and 2-2 are both equal to $\varphi$. With introduction of the variable $\alpha=\varphi-\pi/2$, one obtains $\beta=\pi/2-2\alpha$, where $\beta$ is defined in Fig. 13(a). Denoting the magnitude of $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ by $\lambda_{1}$ and $\lambda_{2}$, respectively, the geometric constraint requires $\tan\beta=(\lambda_{1}-\lambda_{2}\sin\alpha)/(\lambda_{2}\cos\alpha)$. Elimination of the auxiliary variables $\beta$ and $\alpha$ yields
@@
Based on Eq. (6), when $\lambda_{2}/\lambda_{1}\rightarrow 2^{-}$, $\varphi\rightarrow\pi$, corresponding to the two-armed mode limit; when $\lambda_{2}/\lambda_{1}\rightarrow 0^{+}$, $\varphi\rightarrow\pi/2$, indicating a four-armed mode. However, this latter situation is unlikely to occur because of the fact that while $\lambda_{2}$ is finite in general, $\lambda_{1}$, caused by non-contact repulsion, cannot be excessively large. The situation of $\lambda_{2}/\lambda_{1}>2$ is not covered by Eq. (6), either can it be described by the geometrical configuration shown in Fig. 13(a). Actually, it represents the scenario that the non-contact repulsion between bubbles 2-2 and 0 is too small to influence the pattern formation, and the divergence angle is dominated by the pair collision between bubbles 1-1 and 0. Thus, the fixed two-armed mode or disorder mode is expected.
Figure 13: (a) Geometric configuration of bubbles on the free surface (bubble size and position for illustrative purposes only). (b) Simulated bubble trajectories projected onto the horizontal plane after appropriate translation and rotation.
@@
Eq. (7) is valid unless $T$ is too large, in which case disordered, instead of regular, patterns, will be observed. Figure 14(a) shows the averaged $\varphi$ as a function of $T$ for eight simulation cases (black circles) and the experimental data from Ref. [27] (red squares). Notably, with $a$ and $b$ fitted for each data set, the predictions of Eq. (7) (solid lines) have a remarkable consistency with both the numerical and experimental results. The fitted coefficients for the two curves, from left to right, are $a=0.039$ ms-1, $b=0.83$, and $a=0.035$ ms-1, $b=0.90$, respectively.
Figure 14: (a) Divergence angle $\varphi$ versus release period $T$. Red squares denote experimental data extracted from Fig. 1.2(a) in Ref. [27], with $\mathit{Oh}\approx 0.099$ and $\mathit{Bo}=17$. (b) Divergence angle versus dimensionless frequency $\Omega$. The three data sets on the left are experimental results extracted from Fig. 5(a) in Ref. [25]. From left to right, the Ohnesorge and Bond numbers, estimated based on the available information in Ref. [25], are approximately 0.26 and 23, 0.21 and 18, 0.11 and $1$, respectively. In both panels, the present numerical results are denoted by black circles; solid curves represent predictions using Eq. (7) or (8).
@@
where $A=\rho_{l}d^{2}a/\mu_{l}$, and $\Omega=\rho_{l}d^{2}/(\mu_{l}T)$ is the dimensionless bubble release frequency defined in Ref. [25]. The dependence of $\varphi/(2\pi)$ on $\Omega$ is shown in Fig. 14(b). The present numerical results are still denoted by black circles. The other three data sets correspond to experimental data extracted from Fig. 5(a) in Ref. [25]. The results of Eq. (8) are shown as black curves with $A$ and $b$ fitted for each set of data. The values of $A$ and $b$ for the four curves in Fig. 14(b), from left to right, are: 13.3 and 0.87, 16.5 and 0.64, 17.9 and 0.35, 31.1 and 0.83, respectively. Good agreement is observed between the model predictions and all the data sets.
In the following, we consider the possibility of observing different multi-armed patterns in reality. Figure 15(a) shows the number $n$ of arms that is expected to be seen as the release period $T$ varies. This mapping is obtained by calculating the corresponding $\varphi$ at a given $T$ using Eq. (7), then finding $n$ within the range $2\leq n\leq 15$ that best fits Eq. (5) with all possible $m$. The upper bound of 15 is set considering that for very large $n$, the finite-sized bubbles are so crowded that the arms become difficult to distinguish in practice. It can be seen from Fig. 15(a) that each $n$ may correspond to several discontinuous intervals of $T$. The accumulated range of $T$ for a given $n$ is shown in Fig. 15(b). The probability of appearances of $n$-armed mode can thus be estimated based on the height of the bars.
@@ V Summary and conclusions
This work elucidates the formation of regular surface patterns induced by chains of small rising bubbles. Direct numerical simulations reveal how distinct morphologies stem from non-contact bubble repulsion, pair collisions, free-surface deformation, and ambient flow advection. We establish a heuristic model relating the number of pattern arms (represented by the divergence angle $\varphi$) to the bubble release period $T$, showing quantitative agreement with both simulation data and experimental results from the literature. While this heuristic model is not derived from first-principle fluid dynamics, it bridges observed geometric patterns with causal hydrodynamic processes, offering a predictive tool for the self-organized pattern formation in non-coalescing bubble systems. The main predictions of the model include: (1) the existence of stable, discrete surface patterns with fixed arm numbers; (2) a quantitative relationship between the divergence angle of successive bubbles (thus the arm number) and their release period; and (3) preferential occurrence of specific patterns, with three- and five-armed cases being the most probable. It is worthwhile to mention that, the focus of this study is the general relationship between $\varphi$ and $T$. Effects of other parameters–such as bubble diameter, liquid viscosity, and surface tension coefficient–are implicitly included in the coefficients $a$ and $b$ of the proposed model (Eq. (7)). If bubbles have not reached terminal velocity near the free surface, liquid depth becomes an additional parameter that has a quantitative influence on $a$ and $b$; but all essential processes responsible for pattern formation remain unchanged.
@@ ACKNOWLEDGEMENTS
This work was supported by the National Natural Science Foundation of China (Grant No. 12202441) and the Xiaomi Young Scholar Program. D. Li appreciates the advice from Dr. Petr Karnakov on the use of Aphros. The authors also thank Dr. Zibo Ren for his participation in the early stage of this study.
Appendix A Mesh-Independence Assessment
The spiral three-armed pattern in Fig. 4(d) was obtained with a base mesh resolution of $512\times 512\times 192$. To validate this result, an additional simulation for the same case is conducted using a finer mesh ($768\times 768\times 288$). This higher-resolution computation requires more than 20 days if running in parallel on 128 CPU cores. The resulting pattern is presented in Fig. 16(a). Although the pattern rotates in the opposite direction (the direction has no preference, as it is determined by random disturbances during initial bubble interactions), the average divergence angles agree closely (128.8 versus $124.6^{\circ}$), with a difference of less than 3.5%.
Moreover, Fig. 16(b) compares the temporal evolution of the rise velocity for a specific bubble in the bubble chain. The near overlap of the curves from both meshes further justifies the use of the base mesh resolution.
Figure 16: (a) Surface pattern obtained with a mesh resolution of $768\times 768\times 288$ for the three-armed case shown in Fig. 4. (b) Temporal evolution of the bubble rise velocity. The fine- and coarse-mesh results are denoted by dashed and solid lines, respectively.